OpenStudy (anonymous):

The density of solid Cr is 7.15 g/cm3. How many atoms are present per cubic centimeter of Cr?

OpenStudy (anonymous):

\[Density = \frac{ Mass }{ Volume }\]

OpenStudy (anonymous):

You'd want to use the molar mass of Cr ( in units grams/mol ) and then avogadro's number to calculate the number of atoms of Cr per cubic cm. Your units will look as follows\[\frac{ g }{ cm^3 }\times \frac{ mol }{ g }\times \frac{ 6.022 \times 10^{23} atoms}{ mol }\]